# Leveraging Quantum Superposition to Infer the Dynamic Behavior of a Spatial-Temporal Neural Network Signaling Model

### v2

@@ Abstract
The exploration of new problem classes for quantum computation is an active area of research. In this paper, we introduce and solve a novel problem class related to dynamics on large-scale networks relevant to neurobiology and machine learning. Specifically, we ask if a network can sustain inherent dynamic activity beyond some arbitrary observation time or if the activity ceases through quiescence or saturation via an ’epileptic’-like state. We show that this class of problems can be formulated and structured to take advantage of quantum superposition and solved efficiently using the Deutsch–Jozsa and Grover quantum algorithms. To do so, we extend their functionality to address the unique requirements of how input (sub)sets into the algorithms must be mathematically structured while simultaneously constructing the inputs so that measurement outputs can be interpreted as meaningful properties of the network dynamics. This, in turn, allows us to answer the question we pose.
@@ 1 Introduction
Research into quantum computing applications has historically focused on a few well-established topics. These include cryptography, data security, and the simulation of complex physical and chemical systems, including quantum mechanics itself [1, 2, 3, 4, 5, 6].
@@
However, exploring new problem classes that are particularly suitable for quantum computation is an active area of research. This includes both the application of existing quantum algorithms and the development of new ones. For example, Google Quantum and X Prize recently announced a competition to promote the development of practical, real-world quantum computing algorithms and applications (https://www.xprize.org/prizes/qc-apps).
@@
One of the unique challenges of exploring the applicability of quantum computing is that there is no known or obvious process for systematically structuring problems or questions to conform to the mathematical and algorithmic requirements of quantum computation. The structure and solution of each problem are essentially a bespoke pursuit. A well-posed problem needs to leverage quantum properties such as superposition, entanglement, and interference so that measurement outcomes are meaningful and interpretable to the question being addressed. Achieving this requires designing problems that can demonstrate the computational advantages of quantum algorithms while also showing their relevance to practical or theoretical topics and problems of interest.
@@
In this paper, we introduce and solve a novel problem class related to dynamics on large-scale networks relevant to neurobiology and machine learning. Specifically, given a particular spatial (geometric)-temporal neural signaling model, we ask if a network can sustain inherent dynamic activity beyond some arbitrary observation time or if the activity ceases through quiescence or saturation via an ’epileptic’-like state. In a neuroscientific context, this question informs how and when brain networks can encode and represent information given their physical structural and temporal properties. In the context of machine learning, this work relates to understanding activation patterns in artificial neural networks, which have implications for stability, efficiency, and memory.
@@
From a computational perspective, solving this problem classically requires brute-force evaluation of the network’s state by assessing the internal state of each node individually, an approach that is increasingly computationally expensive and eventually prohibitive for very large networks. We show that this problem can be formulated and structured to take advantage of quantum superposition and solved efficiently using the Deutsch–Jozsa and Grover quantum algorithms. To do so, we extend their functionality to address the unique requirements of how input (sub)sets into the algorithms must be mathematically structured while simultaneously constructing the inputs so that measurement outputs can be interpreted as meaningful properties of the network dynamics. This, in turn, allows us to answer the question we pose.
@@
Our intent here is twofold. First, it demonstrates a proof-of-concept application of quantum computing to a network dynamics problem with theoretical significance in neuroscience and machine learning. Second, it provides a systematic framework for thinking about and structuring problems in a way that exploits the advantages of quantum computing, contributing to the broader effort of exploring quantum computing’s applicability to new types of problems and questions.
@@
The logical flow of the paper is as follows: We first introduce the network dynamic model. We then formally define the problem that we aim to answer. We then describe how the quantum algorithms can be adapted and extended to solve the problem efficiently. Specifically, the main intellectual contributions are in the ability to take a sophisticated neurodynamical model and re-frame it in a set-theoretic way that conforms to the structural input requirements of the Deutsch-Jozsa algorithm, such that simultaneously, the measurement outputs are interpretable behaviors of the network dynamics. We then provide formal proofs of complexity advantages and numerical demonstrations of speedups for the quantum versus purely classical approach to solving the problem.
@@ 2.1 Problem Set Up
We aim to address a specifically constructed question regarding the dynamics of neural networks by leveraging the uniquely quantum property of superposition. The specific problem we address is the following: After some period of dynamic evolution, can the dynamics (i.e., activity) of a neural network sustain inherent activity beyond an arbitrary observation time $T_{o}$, or are the dynamics guaranteed to stop, either due to ’epileptic’-like saturation or quiescence?
By inherent dynamic activity, we mean the ability of the neurons (nodes) within the network to continue firing and maintaining the network’s internal dynamics autonomously without requiring an external driving input. The network functions as a self-sustaining system where activity persists due to its internal structure and the interactions between connected nodes.
@@
Much of the technical setup will be in defining and describing the neural dynamic model we will be using. While this is not the only model we could have explored, we chose it because:
@@
    It is computationally tractable and builds directly on prior work.
@@
    It is of relevance to real neurobiological systems. We have previously shown how individual neurons and networks of neurons (connectomes) use the model’s specific theoretical and computational properties as a functional optimization principle.
@@ 2.2 Overview of the Solution
After introducing the model and describing it mathematically, we then show how to map the network dynamics question into a framework suitable for quantum computation. This involves formulating the problem to exploit quantum superposition using the Deutsch-Jozsa and Grover algorithms. We will show that the quantum computations result in interpretable output measurements that enable us to infer whether the network dynamics can sustain inherent activity or cease through epileptic-like saturation or quiescence. By carefully structuring the problem, we will show that a quantum algorithmic solution is computationally more efficient than purely classical methods.
@@
The Deutsch-Jozsa algorithm is a quantum computing algorithm designed to efficiently determine whether a function behaves the same way for all its inputs, outputting either all 0’s or all 1’s (constant), or if the output is an equal mix of 0’s and 1’s (balanced). In our case, we provide the algorithm a list of 0’s and 1’s that encode the activation (firing) patterns representing the network’s dynamics. The algorithm processes this information all at once, leveraging the quantum property of superposition. If the outputs are all 0’s or all 1’s (constant), we interpret the network’s activity pattern as being either in a quiescent state (if all 0’s), where no activity happens, or in an epileptic-like state (if all 1’s), where everything fires at once. If the outputs are not all the same, it implies the network’s dynamics are more complex, with activity in the network having the potential to continue processing information.
@@
However, there’s a challenge: the Deutsch-Jozsa algorithm requires evaluating all possible patterns of activity in the network at once, including patterns that did not actually occur in the network. This would make the algorithm’s output impossible to interpret in terms of the network’s real dynamics. To work around this, we divide the problem into two parts. First, we focus only on the patterns of activity that actually occurred in the network. For the patterns that did not occur, we assign them fixed, known values (either 0 or 1), depending on how we are running the algorithm.
@@
This approach allows us to run the Deutsch-Jozsa algorithm twice in the following way. In the first run, we assign all non-occurring patterns a value of 0. In the second run, we assign them a value of 1. The patterns of activity that actually occurred are already encoded in the collection of observed or measured values of 0’s and 1’s. If the output of the algorithm in the first run (where non-occurring patterns were assigned 0’s) is all 0’s, this implies that the activity patterns in the network were also all 0’s, meaning the network was in a quiescent state. Similarly, if the output of the second run (where non-occurring patterns were assigned 1’s) is all 1’s, it implies that the network was in an epileptic-like state. If the outputs are not constant (i.e., not all 0’s or not all 1’s) in either case, it implies that the activity patterns in the network were a mix of 0’s and 1’s, and we interpret this as the network having the potential to sustain dynamic activity.
@@
By interpreting the outputs of these two runs, we can derive the state of the activity pattern of the network and determine if the activity can continue or if it stops. The quantum computational advantage lies in the fact that we do not need to evaluate each pattern individually, one at a time. Instead, the Deutsch-Jozsa algorithm evaluates the entire set of patterns simultaneously in superposition, saving an enormous amount of computational effort. The larger the network being interrogated, the greater the computational speedup and resource savings achieved through this approach.
@@ 3 Refractory and Signal Latency Dynamics in Geometric Neural Networks
The network dynamic model we will use operates on structural geometric networks. It is effectively a sophisticated geometric extension of the classical integrate-and-fire model. This model takes into account the neurobiologically canonical processes of spatial and temporal summation of discrete signaling events (action potentials) incident on nodes (neurons) within a geometric network. We define a geometric network as a physically constructible structural network where the edges between neurons are convoluted paths with physical distances. In a neurobiological context, this corresponds to the path lengths of axons and dendrites that connect chemical synapses between neurons, making the model anatomically accurate. The inclusion of network geometry and finite signaling speeds (conduction velocities) of action potentials naturally introduces signaling latencies.
@@
In the remainder of this section, we provide a self-contained technical description of the model. Readers interested in its full development, associated proofs, and theoretical and experimental results are referred to the references listed below.
@@ 3.3.1 Geometric network construction
We write $H_{j}(v_{i})$ to denote the set of all vertices $v_{i}$ in $H_{j}$, and $H_{j}[v_{i}]$ to refer to a specific $v_{i}\in H_{j}(v_{i})$. Signals traveling on the edges $e_{ij}$ propagate at finite speeds $s_{ij}$, where $0<s_{ij}<\infty$. As a first approximation, we assume that a single arriving discrete signal has the potential to activate $v_{j}$. To indicate causality, we use the notation $H_{j}[v_{i}]\leadsto v_{j}$ to signify that vertex $v_{i}\in H_{j}(v_{i})$ causally activates $v_{j}$, meaning that $v_{i}$ is the winning vertex whose signal successfully activates $v_{j}$. We first treat and analyze this restricted (single activation) case and 3.3.3) and then generalize the model to a latency-determined spatial-temporal version of a perceptron that activates due to the fractional summation of many arriving discrete signals (section 3.5.3).
@@ 3.3.2 Refractory period and signal dynamics
Consider a vertex $v_{i}$ with a directed edge $e_{ij}$ connecting to a vertex $v_{j}$. For $v_{i}$ to signal $v_{j}$, a discrete signal must propagate along $e_{ij}$ at a finite speed $s_{ij}$. While $s_{ij}$ could be constant for all edges, this is not required in the general case. Similarly, while nodes in the network may share the same refractory period (i.e., $R_{j}=R\;\forall v_{j}\in V$), the framework accommodates node-specific values of $R_{j}$.

### v4

@@ Abstract
The exploration of new problem classes for quantum computation is an active area of research. In this paper, we introduce and solve a novel problem class related to dynamics on large-scale networks relevant to neurobiology and machine learning. Specifically, we ask if a network can sustain inherent dynamic activity beyond some arbitrary observation time or if the activity ceases through quiescence or saturation via an ’epileptic’-like state. We show that this class of problems can be formulated and structured to take advantage of quantum superposition and solved efficiently using a coupled workflow between the Grover and Deutsch–Jozsa quantum algorithms. To do so, we extend their functionality to address the unique requirements of how input (sub)sets into the algorithms must be mathematically structured while simultaneously constructing the inputs so that measurement outputs can be interpreted as meaningful properties of the network dynamics. This, in turn, allows us to answer the question we pose.
@@ 1 Introduction
Research into applications of quantum computing has historically focused on a few well-established topics. These include cryptography, data security, and the simulation of complex physical and chemical systems, including quantum mechanics itself [1, 2, 3, 4, 5, 6].
@@
However, exploring new problem classes that are particularly suitable for quantum computation is an active area of research. This includes both the use of existing quantum algorithms and the development of new ones. For example, Google Quantum and X Prize recently announced a competition to promote the development of practical, real-world quantum computing algorithms and applications (https://www.xprize.org/prizes/qc-apps).
@@
One of the unique challenges of exploring the applicability of quantum computing is that there is no known or obvious process for systematically structuring problems or questions to conform to the mathematical and algorithmic requirements of quantum computation. The structure and solution of each problem are essentially a bespoke pursuit. A well-posed problem needs to leverage quantum properties such as superposition, entanglement, and interference so that measurement outcomes are meaningful and interpretable to the question being addressed. Achieving this requires structuring problems that can demonstrate the computational advantages of quantum algorithms while also showing their relevance to practical or theoretical topics and problems of interest.
@@
In this paper, we introduce and solve a novel problem class related to dynamics on networks, broadly relevant to neurobiology and machine learning. We demonstrate that our solution is theoretically provably solvable more efficiently using two canonical quantum algorithms and show quantum simulation results for a small network toy example implementation of the problem. The question we address is the following: We ask if a network, consisting of nodes (neurons) whose dynamics are governed by a particular neural signaling model, can sustain inherent dynamic activity beyond some arbitrary observation time or if the activity ceases through quiescence or saturation via an ’epileptic’-like state. In a neuroscientific context, this question informs how and when brain networks can encode and represent information given their physical structural and temporal properties. In the context of machine learning, this work relates to understanding activation patterns in artificial neural networks, which have implications for stability, efficiency, and memory.
@@
From a computational perspective, solving this problem classically requires brute-force evaluation of the network’s global state by assessing the internal state of each node individually, an approach that is increasingly computationally expensive and eventually prohibitive for very large networks. We show that this problem can be formulated and structured to take advantage of quantum superposition and solved efficiently in a workflow that leverages in a connected way the Grover and Deutsch–Jozsa quantum algorithms. To achieve this, we extend their functionality so that the framework we develop conforms to the unique requirements of how input (sub)sets into the algorithms must be mathematically structured, while simultaneously constructing the inputs to ensure that measurement outputs can be interpreted as meaningful properties of the network dynamics. This, in turn, allows us to answer the question we pose.
@@
Our intent here is twofold. First, it demonstrates a proof-of-concept application of quantum computing to a network dynamics problem with theoretical significance in neuroscience and machine learning. Second, it provides an approach for thinking about and structuring network dynamics problems in a way that exploits the advantages of quantum computing, contributing to the broader effort of exploring quantum computing’s applicability to new types of problems and questions.
@@
The logical flow of the paper is as follows: We first introduce the network dynamic model. We then formally define the problem that we aim to answer. We then describe how the quantum algorithms can be adapted and extended to solve the problem efficiently. We then provide formal proofs of complexity advantages and numerical demonstrations of speedups for the quantum versus purely classical approach to solving the problem. We conclude by constructing and illustrating a set of chained quantum circuits and a logical workflow, which we use to numerically simulate and solve representative toy examples of the problem by implementing our theoretical approach.
The main intellectual contribution of this work is the ability to take a neurodynamical model and reframe it in a set-theoretic manner that conforms to the mathematical and computational constraints of the quantum algorithms, while simultaneously rendering the measurement outputs as interpretable behaviors of the network dynamics.
@@ 2.1 Problem Set Up
Our goal is to address a specific question regarding the dynamics of neural networks by leveraging the uniquely quantum property of superposition. The specific problem we address is the following: After some period of dynamic evolution, can the dynamics (i.e., activity) of a neural network sustain inherent activity beyond an arbitrary observation time $T_{o}$, or are the dynamics guaranteed to stop, either due to ’epileptic’-like saturation or quiescence?
By inherent dynamic activity, we mean the ability of the neurons (nodes) that make up the network to continue firing and maintaining the network’s internal dynamics autonomously without requiring an external driving input. The network functions as a self-sustaining system where activity persists due to its internal structure and the interactions between connected nodes.
@@
Much of the technical setup will be in defining and describing the neural dynamic model. While the model we introduce is certainly not the only model we could have explored, we chose it because:
@@
    It is computationally tractable and builds directly on prior work, but resource-intensive and scales with network size, i.e., the number of nodes and edges.
@@
    It is of relevance to real neurobiological systems. We have previously shown how individual neurons and networks of neurons (connectomes) use the model’s specific theoretical and computational properties as a functional optimization principle. See section 3.2 below.
@@
    It is of relevance to machine learning in that we have previously shown how the model can be used to carry out a type of data (vector) embedding and encoding whereby the relationships between encoded data are inherently a component of measurable distances between vectors in the embedding space [9, 10].
  5. 5.
@@
  6. 6.
    And lastly, taken all together, the model can set the stage for follow on exploratory work as a test case and benchmark for related quantum computing problem classes and models.
@@ 2.2 Overview of the Solution
After introducing the model and describing it mathematically, we show how to map the network dynamics question into a framework suitable for quantum computation. This involves formulating the problem to exploit quantum superposition using the Grover and Deutsch-Jozsa algorithms. We will show that the output measurements from the quantum computation result in interpretable results that allow us to infer whether the network dynamics can sustain inherent activity or cease through epileptic-like saturation or quiescence. By carefully structuring the problem, we will show that a quantum algorithmic solution is computationally more efficient than purely classical methods.
@@
The Deutsch-Jozsa algorithm is a quantum computing algorithm designed to efficiently determine whether a function behaves the same way for all its inputs, outputting either all 0’s or all 1’s (constant), or if the output is an equal mix of 0’s and 1’s (balanced). In our case, we provide the algorithm a list of 0’s and 1’s that encode the activation (firing) patterns representing the network’s dynamics. The algorithm processes this information all at once, leveraging the quantum property of superposition. If the outputs are all 0’s or all 1’s (constant), we interpret the network’s activity pattern as being either in a quiescent state, where no activity happens, or in an epileptic-like state, where everything fires at once. If the outputs are not all the same, it implies the network’s dynamics are more complex, with activity in the network having the potential to continue processing information.
@@
However, there’s a challenge: the Deutsch-Jozsa algorithm requires evaluating all possible patterns of activity in the network at once, including patterns that did not actually occur in the network. This would make the algorithm’s output impossible to interpret in terms of the network’s real dynamics. To work around this, we divide the problem into two parts. Critically, we use Grover’s algorithm to identify and amplify only the patterns of activity that actually occurred in the network. For the patterns that did not occur, we assign them fixed, known values (either 0 or 1), depending on how we are running the algorithm.
@@
This approach allows us to run the Deutsch-Jozsa algorithm twice in the following way. In the first run, we assign all non-occurring patterns a value of 0. In the second run, we assign them a value of 1. The patterns of activity that actually occurred are already encoded in the collection of observed or measured values of 0’s and 1’s. If the output of the algorithm in the first run (where non-occurring patterns were assigned 0’s) is all 1’s, this implies that the activity patterns in the network were also all 0’s, meaning the network was in a quiescent state. (Note that the algorithm ‘flips’ the all 0 input to all 1 outputs due to technical reasons of how it operates and computes.) Similarly, if the output of the second run (where non-occurring patterns were assigned 1’s) is all 0’s, it implies that the network was in an epileptic-like state. If the outputs are not constant (i.e., not all 0’s or not all 1’s) in both cases, it implies that the activity patterns in the network were a mix of 0’s and 1’s, and we interpret this as the network having the potential to sustain dynamic activity. This is a unique non-standard application of Deutsch-Jozsa that leverages the problem’s structure and setup (see Section 2.3 immediately below).
@@
By interpreting the outputs of these two runs, we can infer the state of the activity pattern of the network and determine if the activity can continue or stop. The quantum computational advantage lies in the fact that we do not need to evaluate each pattern individually, one at a time. Instead, the combined use of Grover, linked with two runs of Deutsch-Jozsa, evaluates the entire set of patterns simultaneously in superposition, saving an increasing amount of computational effort as the network size increases. The larger the network being interrogated, the greater the computational speedup and resource savings this approach achieves.
2.3 Beyond classical oracle limitations: Why the standard criticisms do not apply
In this section, we attempt to address upfront how structured Grover amplification supports the practical use of Deutsch-Jozsa, given the specific construction of the problem we introduce, and why the standard criticisms and limitations of Deutsch-Jozsa in practical settings do not apply. Our goal here is to convince readers, or at least intrigue them enough, that the typical arguments against any Deutsch-Jozsa quantum advantage in a real-world problem are not valid, in the hope that they will continue reading. We assert that, in at least some instances, Deutsch-Jozsa, when combined with Grover and possibly other algorithms as part of a broader solution workflow, does confer a computational speedup and advantage.
In the standard textbook case, Deutsch-Jozsa is analyzed under a very specific and constrained set of conditions. One is given a black-box oracle function that evaluates to the binary output
$f:\{0,1\}^{n}\rightarrow\{0,1\}$
Built into this construction is an explicit promise that the oracle function is either constant, i.e., all outputs are either 00 or $1$, or balanced, i.e., exactly half the outputs are 00 and exactly half are $1$. The objective is to determine with certainty and no error whether $f$ is constant or balanced. Under these conditions, Deutsch-Jozsa is guaranteed to arrive at a solution in one attempt.
In most real-world conditions, the strict textbook promise of a function being exactly constant or exactly balanced does not apply. It simply is not a condition of relevance to the structure of real-world systems and the physical questions asked of them. Functions may be almost balanced or almost constant, but noise and approximation conditions render the exactness constraint unrealistic. The problem is that if one relaxes this constraint and allows even a small error in the acceptable evaluation of the function, then a classical randomized algorithm operating on a small subset of samples will quickly be able to distinguish between almost balanced or almost constant conditions with high probability every time. There is no advantage to a quantum algorithm like Deutsch-Jozsa, and therefore, no need for quantum computing.
However, how we structure the question we ask in this paper, and the approach we use to solve it, deviate from the textbook usage of the Deutsch-Jozsa algorithm while respecting the strict exactness constraints when we do make use of it. We do not rely on Deutsch-Jozsa to directly answer the question about network dynamics; we use it strictly as a final ’checking’ substep in the workflow on a simple function that meets the textbook criteria, but which is interpretable to us in the broader context of how the problem was intentionally set up.
Specifically, we are not evaluating a random function with noise or deviations from the constant or balanced exactness constraints. The function we evaluate is an actualized set of structured spiking events that, by construction, is (or is not) always exactly constant. We use Grover’s algorithm first to efficiently identify and amplify only the subset of states that were actually realized in a specific evolution of the network dynamics, using an explicitly defined thresholding function. All non-actualized states are uniformly set to either 0’s or 1’s (in two different back-to-back runs). The Grover quantum pre-processing step is critical, and classical random sampling cannot provably be as efficient. This is a source of (quadratic) quantum speedup in the workflow.
Because we construct the overall function ourselves, we can ensure an exactly constant scenario, i.e., all states are 0 (if the network is quiescent) or 1 (if the network is epileptic). If both runs deviate from exact consistency, it can only be due to a mixture of 0’s and 1’s, which we can neurophysiologically interpret as a potential for continued dynamics. In other words, we guarantee the exactness constraint by construction, and can ’check’ it provably faster using canonical quantum algorithms versus classically necessarily ’sequentially’ checking the state of each node (neuron) in the network.
In this workflow, Deutsch-Jozsa is merely a final validation step after marking states using Grover’s algorithm. It operates on a constructively exact constant function. But it is only a step in a sequence of steps that allow us to interpret the final outputs of the workflow in a meaningful neurophysiological context. What we do not do is attempt to use Deutsch-Jozsa on an approximate function that relaxes the exactness constraints as the end-all and be-all.
@@ 3 Refractory and Signal Latency Dynamics in Geometric Neural Networks
The network dynamic model we will use operates on structural geometric networks. From a technical perspective, it is effectively a geometric extension of classical neuronal models. This model takes into account the neurobiologically canonical processes of spatial and temporal summation of discrete signaling events (action potentials) incident on nodes (neurons) within a geometric network. We define a geometric network as a physically constructible structural network where the edges between neurons are convoluted paths with physical distances. In a neurobiological context, this corresponds to the path lengths of axons and dendrites that connect chemical synapses between neurons, making the model anatomically accurate. The inclusion of network geometry and finite signaling speeds (conduction velocities) of action potentials naturally introduces signaling latencies.
@@
In the remainder of this section, we provide a self-contained technical description of the model. Readers interested in its full development, associated proofs, and theoretical and experimental results are referred to the references cited below.
@@ 3.3.1 Geometric network construction
We write $H_{j}(v_{i})$ to denote the set of all vertices $v_{i}$ in $H_{j}$, and $H_{j}[v_{i}]$ to refer to a specific $v_{i}\in H_{j}(v_{i})$. Signals traveling on the edges $e_{ij}$ propagate at finite speeds $s_{ij}$, where $0<s_{ij}<\infty$. As a first approximation, we assume that a single arriving discrete signal has the potential to activate $v_{j}$. To indicate causality, we use the notation $H_{j}[v_{i}]\leadsto v_{j}$ to signify that vertex $v_{i}\in H_{j}(v_{i})$ causally activates $v_{j}$, meaning that $v_{i}$ is the winning vertex whose signal successfully activates $v_{j}$. We first treat and analyze this restricted (single activation) case and then generalize the model to a latency-dependent spatial-temporal version of a perceptron that activates due to the fractional summation of many arriving discrete signals (section 3.5.3).
@@ 3.3.2 Refractory period and signal dynamics
Consider a vertex $v_{i}$ with a directed edge $e_{ij}$ connecting to a vertex $v_{j}$. For $v_{i}$ to signal $v_{j}$, a discrete signal must propagate along $e_{ij}$ at a finite speed $s_{ij}$. While $s_{ij}$ could be a constant for all edges, this is not required in the general case. Similarly, while nodes in the network may share the same refractory period (i.e., $R_{j}=R\;\forall v_{j}\in V$), the framework can accommodate node-specific values of $R_{j}$.